Suppose x, y, z is a geometric sequence with common ratio r and x≠ y. If x, 2y, 3z is an arithmetic sequence, then what is r^2?
Answer by ikleyn(52803)

.
I will only give you a HINT, leaving the solution to you.


HINT 1.  Since x, y ans z form Geom.progression, y = rx and z = r^2*x.       (1)


HINT 2.  Since x, 2y and 3z form Arithm.progression, 2y-x = 3z - 2y.         (2)


HINT 3.  Having this, substitute (1) to (2).  You will get an equation


2(r*x) -x = 3(r^2*x) - 2*(r*x).


Cancel x in both sides and solve for "r".

Happy solving !!